OpenStudy (anonymous):

write an equation of the line containing the given points & perpendicular to the given line. (0,5) ; 8x+5y=3

OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (anonymous):

(-8x+3)/5=y

OpenStudy (anonymous):

y=-8/5x +3/5

OpenStudy (anonymous):

any questions on what I did so far?

OpenStudy (anonymous):

can you show me how you got that?

OpenStudy (anonymous):

your equation was 8x+5y=3 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I was equation of y in terms of x

OpenStudy (anonymous):

wait is this the answer to my question?

OpenStudy (anonymous):

So I subtracted 8x from both side I got 5y=-8x+3

OpenStudy (anonymous):

any questions so far djm?

OpenStudy (anonymous):

so this isnt the answer to my question?

OpenStudy (anonymous):

don't quite understand your statement of equation of y in terms of x

OpenStudy (anonymous):

we are describing y in terms of x another way to explain would be what do you have to do to x so it equals y

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

ya that is a little better

OpenStudy (anonymous):

so getting back to our equation 5y=-8x+3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we want y by itself so we divide by 5 BOTH SIDES.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=-8/5x+3/5

OpenStudy (anonymous):

any questions so far?

OpenStudy (anonymous):

no good so far

OpenStudy (anonymous):

Now we have to go back to the original question

OpenStudy (anonymous):

ok sorry had to refresh again

OpenStudy (anonymous):

i need help please T_T no one wants to help

OpenStudy (anonymous):

"perpendicular to the given line" is the keyword from your question

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

PLEASE PLEASE! IM SORRY! BUT I NEED HELP,T___T

OpenStudy (anonymous):

parallel lines have same slope perpendicular lines have negative reciprocal slope

OpenStudy (anonymous):

right

OpenStudy (anonymous):

going back to our equation y=-8/5x+3/5

OpenStudy (anonymous):

so noes going to help me!

OpenStudy (anonymous):

>:( fine, feel my wrath of rage

OpenStudy (anonymous):

slope is -8/5 opposite reciprocal is 5/8 point slope form is y - y1=m(x-x1) from question (y-5)=5/8(x-0)

OpenStudy (anonymous):

y-5=5/8(x-0) that is your equation any question

OpenStudy (anonymous):

ok that is where I was having problems was with the last formula u used. thank you very much for taking the time to explain it to me.